The walkway system has an efficiency of 78%. Calculate the power input to the system when

15 passengers of average mass 72 kg are standing on the walkway.


1
Expert's answer
2021-10-18T15:06:19-0400

The weight of the passengers is



"W = mg"

where "m = 1080kg" is they mass, and "g = 9.8N\/kg" is the gravitational acceleration. The component of the weight that acts along the walkway is:



"W_a = W\sin \theta"

where "\theta = 30\degree".

By definition, the power is:



"P = Wv"

where "v = 0.51m\/s" is the speed of motion. Thus, obtain:



"P = mgv\sin\theta\\
P = 1080\cdot 9.8\cdot 0.51\cdot \sin 30\degree \approx 2698.92W"


Then, given the efficiency of the walkway, the power input has to be equal to the required power input (2698.92 W) divided by the efficiency (78% = 0.78):

"\text{efficiency} = \frac{\text{Power delivered}}{\text{Power input}}
\\ \implies \text{Power input} = \frac{\text{Power delivered}}{\text{efficiency}}= \frac{2698.92 \text{ W}}{0.78}
\\ \implies \text{Power input} = 3460.15 \text{ W}"

Answer. 3460.15 W.